OpenStudy (anonymous):

Find x if y=32. y=(x/32)X.125

OpenStudy (anonymous):

Okay, so for this you need to substitute. Do you know what that means?

OpenStudy (anonymous):

yes, but im not sure exactly how to

OpenStudy (anonymous):

Okay, so it says that y = 32. So for why in you equation put 32 in for where y is. :)

OpenStudy (anonymous):

32 = (x/32) * 125. (It is times 125, right?

OpenStudy (anonymous):

so it would be 32-(x/32)X.125 yes, it is times

OpenStudy (anonymous):

correct. what is with the X. though? is that times?

OpenStudy (anonymous):

sorrry, = not - ya, the X is the times

OpenStudy (anonymous):

okay. we will change the X. to be * to make it where we don't confuse the X and the x.

OpenStudy (anonymous):

ok, that makes sense

OpenStudy (anonymous):

So now, you need to get the 32 out of the fraction bar. do you know how to do that?

OpenStudy (anonymous):

wouldnt i have to multiply both sides by 32?

OpenStudy (anonymous):

we need to get the 32 out so x will be by itself.

OpenStudy (anonymous):

correct!

OpenStudy (anonymous):

so what is 32 * 32?

OpenStudy (anonymous):

(you can use a calculator)

OpenStudy (anonymous):

1024

OpenStudy (anonymous):

okay, so now we have 1,024 = x * 125 (or 125x, so it is actually 1,024 = 125x.) We will change this to make it easier. It is now 125x = 1024. what now?

OpenStudy (anonymous):

well, its .125 (1/8)

OpenStudy (anonymous):

I don't know exactly what you did, but you divide the 125. So you have x = 1024/125. Us a calculator.

OpenStudy (anonymous):